Suppose an opaque jar contains 5 red marbles and 10 green marbles. I experiment picking 2 marbles from the jar without replacing the first one.
1. What is the probability of getting a green marble first and a red marble second?
2. What is the probability of getting a green marble and a red marble? How is this different from question #1?
For the 1st question I did this: (5x10)/(15x14)= 50/210 = 5/21 (am pretty sure that is right)
For the 2nd question I am stumped...I thought it would be the same answer but something keeps nagging at me that I am missing something. I thought about this: (5)squared +(10)squared/(15x14) = 125/210.595 or 59.5% I really don't understand how the questions are that different..or at least I know they are asking 2 different probable outcomes but don't know how to explain the difference. Can you please help me? Thanks!
Glennys




Answer by stanbon(57347)

Suppose an opaque jar contains 5 red marbles and 10 green marbles. I experiment picking 2 marbles from the jar without replacing the first one.
1. What is the probability of getting a green marble first and a red marble second?
Ans: (10/15)(5/14) = 50/210 = 5/21
----
2. What is the probability of getting a green marble and a red marble? How is this different from question #1?
# of ways to succeed: 5*10 = 50
# of random pairs: 15C2 = 105
---
P(red AND green) = 50/105 = 10/21

=============
Cheers,
Stan H.
=============